The ten-year bond yields 5.5% and has a coupon of 7.6%. If this yield to maturity remains unchanged, what will be its price one year hence? Assume annual coupon payments and a face value of $100
Answer #1

Answer:

Price one year hence:

Balance years to maturity = 10 - 1 = 9 years

Face value = $100

Annual coupon amount = $100 * 7.6% = $7.60

Yield to maturity remains same at = 5.5%

To calculate Price we will use PV function of excel:

PV ( rate, nper, pmt, fv, type)

PV (5.5%, 9, -7.60, -100, 0)

=$114.60

Price one year hence = $114.60